julga,vuyani and khusela are sharing an amount of R176 in the ratio 4:5:7 .calculate how much money each of them wil receive.
Answer by nerdybill(7384)

julga,vuyani and khusela are sharing an amount of R176 in the ratio 4:5:7 .calculate how much money each of them wil receive.
.
Total parts:
4+5+7 = 16
.
Each part is worth:
R176/16 = R11
.
julga: 4*11 = R44
vuyani: 5*11 = R55
khusela: 7*11 = R77